OpenStudy (anonymous):

proove that 1\2 +1\4 +1\8...=1

OpenStudy (anonymous):

Logicly it's obvious, mathematicly please

OpenStudy (anonymous):

putting all of the fractions to a common denominator will help to prove this equation. after some experiments, you would notice that all of those fractions share a common denominator of 8. Hence, the equation can be rewritten as: 4/8 + 2/8 + 1/8 = 8/8. now it is simply a matter of seeing whether this is correct. mathematically, 4+2+1=7, and not 8. Therefore, you can disprove the original equation hope it helped, cheers

OpenStudy (anonymous):

you forget the 3 little dots this is easy, i meant if it went up to 1\2^(infinit)

OpenStudy (anonymous):